OpenStudy (anonymous):

log base 8 (0.25) compute the log

OpenStudy (anonymous):

-2/3

OpenStudy (anonymous):

You can think of it as => 8^(x) = 0.25? What does x need to be.

OpenStudy (dumbcow):

you could also use change of base formula =ln(0.25)/ln(8)

OpenStudy (anonymous):

when i tried to solve it I broke it down to 8^y = 25/100. simplifed and got 8^y = 1/4. 2^3= 1/2^2. the i got stuck

OpenStudy (anonymous):

i know change of base works but i must show work

OpenStudy (anonymous):

8^x = 1/4 we know x has to be -, 8^(2/3) = 4

OpenStudy (anonymous):

Or if you're like me and you can never remember the change of base formula... Let \(k=log_8 (.25)\) \[\implies 8^k = .25\] \[\implies k(ln8) = ln (.25)\] \[\implies k = \frac{ln(.25)}{ln8}\] \[\implies log_8(.25) = \frac{ln(.25)}{ln8}\]

OpenStudy (anonymous):

8^2 = 64 third root of 64 = 4. The way you are showing you work if you can reason that.

OpenStudy (anonymous):

Ok I see